# Partial Semigroupoid Actions on Sets

### v1

@@ 3. The universal globalization
(i) We first verify (G1). If $s,t\in S$ are such that $S^{s}=S^{t}\neq\emptyset$ and $[a,x]\in{{}_{s}E}$, then $[a,x]=[r,y]$ for some $r\in S^{s}$, or $[a,x]=[\delta,y]$ for some $y\in{{}_{\rho_{s}}X}$. In the first case we have $r\in S^{s}=S^{t}$ and, hence, $[a,x]\in{{}_{t}E}$. In the second, $S^{s}=S^{t}\neq\emptyset$ implies $\rho_{s}=\rho_{t}$, hence $y\in{{}_{\rho_{t}}X}$. So $[a,x]=[\delta,y]\in{{}_{t}E}$. This shows ${{}_{s}E}\subseteq{{}_{t}E}$. Since $S^{s}=S^{t}\neq\emptyset$ is a symmetric condition, the same argument proves ${{}_{s}E}={{}_{t}E}$. For the remaining of the proof for (i) we uses Lemma 2.10(ii). That is, for $(s,t)\in S^{(2)}$ we must show ${{}_{t}E}={{}_{st}E}$, $\beta_{t}({{}_{t}E})\subseteq{{}_{s}E}$ and $\beta_{s}\circ\beta_{t}=\beta_{st}$.
If $[a,x]\in{{}_{t}E}$, then $\beta_{t}([a,x])=[tr,y]$ for some $r\in S^{t}$ or $\beta_{t}([a,x])=[t,y]$. Since $t,tr\in S^{s}$, in both cases we have $\beta_{t}([a,x])\in{{}_{s}E}$. So $E_{t}\subseteq{{}_{s}E}$. Since $(s,t)\in S^{(2)}$ implies $S^{t}=S^{st}$ and, by Lemma 3.2(i), $\rho_{t}=\rho_{st}$, we have
@@
(ii) We first verify that $(\delta,\beta,E)$ is a globalization for $(\alpha,X)$. Let $x\in{{}_{s}X}$ and notice that ${{}_{s}X}\subseteq{{}_{\rho_{s}}X}$. Then $(\delta,x)\in{{}_{s}D}$ and, therefore, $\delta(x)=[\delta,x]\in{{}_{s}E}$. Furthermore, for such $x$ we have
@@
So $\delta$ is a morphism of partial actions. It follows from Lemma 3.4(v) that $\delta$ is injective. Therefore, to $\delta$ to be an embedding, it remains to prove that ${{}_{s}X}=\{x\in X\colon\delta(x)\in{{}_{s}E},\beta_{s}(\delta(x))\in\delta(X)\}$. Suppose that $x\in X$ is such that $\delta(x)\in{{}_{s}E}$ and $\beta_{s}(\delta(x))\in\delta(X)$. From $\delta(x)\in{{}_{s}E}$ we have two possible cases:
@@
    $[\delta,x]=[\delta,y]$ for some $y\in{{}_{\rho_{s}}X}$. In this case, Lemma 3.4(v) implies $x=y$. Therefore $\beta_{s}(\delta(x))=[s,x]$. Now, if $[s,x]\in\delta(X)$, then there is $z\in X$ such that $[s,x]=[\delta,z]$. As any chain from $(\delta,z)$ to $(s,x)$ must use relation (1), it follows from Lemma 3.4(ii) that $x\in{{}_{s}X}$.
@@
    $[\delta,x]=[t,y]$, for some $t\in S^{s}$. Similar to the previous case, we obtain from Lemma 3.4(ii) that $y\in{{}_{t}X}$ and $\alpha_{t}(y)=x$. Now, if $\beta_{s}(\delta(x))=[st,y]\in\delta(X)$, then there is $z\in X$ such that $[st,y]=[\delta,z]$. Again, it must be $y\in{{}_{st}X}$. Therefore
@@
This shows that $\{x\in X\colon\delta(x)\in{{}_{s}E},\beta_{s}(\delta(x))\in\delta(X)\}\subseteq{{}_{s}X}$. The reverse inclusion follows from $\delta$ being a morphism of partial actions. This concludes that $(\delta,\beta,E)$ is aglobalization for $(\alpha,X)$.
Now we verify that $(\delta,\beta,E)$ is a universal globalization. For it, let $(\theta,Y)$ be a global action and $\varphi\colon(\alpha,X)\to(\theta,Y)$ be a morphism. Define $\Phi\colon D\to Y$ as
@@
For $\Phi$ to be well defined, we need to verify that $\varphi(x)\in{{}_{s}Y}$ whenever $(s,x)\in D$. If $(s,x)\in D$, then $x\in{{}_{\rho_{s}}X}$ and, hence, we have two possible cases:
@@
    $x\in{{}_{t}X}$ for some $t\in S$ such that $sRt$. In this case, since $\varphi$ is a morphism of partial actions we have $\varphi(x)\in{{}_{t}Y}$. As $(\theta,Y)$ is a global action, it follows from Lemma 3.2(iii) that $\varphi(x)\in{{}_{t}Y}={{}_{s}Y}$.
@@
So the function $\Phi$ is well defined. Furthermore, $\Phi$ is compatible with the relation $\sim$ on $D$. For if $(s,x)\sim(\delta,y)$, then $x\in{{}_{s}X}$ and $y=\alpha_{s}(x)$. Therefore
@@
While if $(s,x)\sim(t,y)$, then there is $u\in S^{t}$ such that $s=tu$, $x\in{{}_{u}X}$ and $\alpha_{u}(x)=y$. Therefore
@@
The case $t=su$ is analogous. As $\Phi$ is compatible with $\sim$ and, consequently, with $\approx$, it induces a function $\Phi\colon E\to Y$, given by $\Phi([a,x])=\Phi(a,x)$. The function $\Phi$ is a morphism of partial actions and satisfy $\Phi\circ\delta=\delta$. For if $[t,x]\in{{}_{s}E}$, where $t\in S^{s}$, then
@@
while if $[\delta,x]\in{{}_{s}E}$, then
$(\Phi\circ\beta_{s})([\rho_{s},x])=(\theta_{s}\circ\varphi)(x)=(\theta_{s}\circ\Phi)([\delta,x]).$
So $\Phi$ is a morphism of partial action. For every $x\in X$ we have $(\Phi\circ\delta)(x)=\Phi([\delta,x])=\phi(x)$. For the uniqueness of $\Phi$, suppose that $\Psi\colon E\to Y$ is a morphism such that $\Psi\circ\delta=\varphi=\Phi\circ\delta$. For each $[\delta,x]\in E$ we have $\Psi([\delta,x])=(\Psi\circ\delta)(x)=(\Phi\circ\delta)(x)=\Phi([\delta,x])$. Now, for each $[s,x]\in E$ we have $x\in{{}_{\rho_{s}}X}$. Therefore $[s,x]=\beta_{s}([\delta,x])$ and, hence,
@@
That is, $\Psi\equiv\Phi$. Therefore $(\delta,\beta,E)$ is a universal globalization for $(\alpha,X)$. ∎
To end this section, we will prove that the universal globalization of a semigroupoid partial action is non-degenerate if and only if the partial action is non-degenerate. Let $(\alpha,X)$ be a partial semigroupoid action. We denote
@@
That is, $X_{1}$ and $X_{0}$ are, respectively, the non-degenerate and the degenerate parts of $X$.
@@
Let $(\alpha,X)$ be a semigroupoid partial action of $S$, and let $(\delta,\beta,E)$ be its universal globalization, as stated in Theorem 3.7. Then $\delta(X_{0})=E_{0}$. Consequently, $(\alpha,X)$ is non-degenerate if and only if $(\beta,E)$ is non-degenerate.
@@
First, we will prove the inclusion $\delta(X_{0})\subseteq E_{0}$. Let $x\in X_{0}$ and $\delta(x)=[\delta,x]\in E$. Suppose that $\delta(x)\in{{}_{s}E}$, for some $s\in S$, and recall that ${{}_{s}E}={{}_{s}D}/\approx$, where
@@
Since $x\in X_{0}$, it follows that $x\notin{{}_{\rho_{s}}X}$. Therefore, $(\delta,x)\simeq(t,y)$ for some $t\in S^{s}$. This implies that there exists a chain of the form
@@
for some $r\in S$. However, $(\delta,x)\sim(r,z)$ means that $x=\alpha_{r}(z)\in X_{r}\subseteq X_{1}$, a contradiction. Now, suppose that $\delta(x)\in E_{s}$, for some $s\in S$. Then, it must be
$[\delta,x]=\beta_{s}([\delta,y])=[s,y]\quad\text{or}\quad[\delta,x]=\beta_{s}([t,y])=[st,y].$
Both cases lead to $(\delta,x)\approx(r,y)$, for some $r\in S$, which, again, result in a contradiction. Therefore
$\delta(x)\notin\cup_{s\in S}({{}_{s}E}\cup E_{s})=E_{1}\implies\delta(x)\in E\setminus E_{1}=E_{0}.$
This concludes that $\delta(X_{0})\subseteq E_{0}$.
To prove the inclusion $E_{0}\subseteq\delta(X_{0})$, we will first consider the case $[a,y]\in E_{0}\cap\delta(X)$. That is, $[a,y]=[\delta,x]$, for some $x\in X$. Suppose that $x\in X_{1}$. Then there is $s\in S$ such that $x\in{{}_{s}X}\cup X_{s}$. Since $\delta\colon(\alpha,X)\to(\beta,E)$ is a morphism of partial actions, if $x\in{{}_{s}X}$, then
$[a,y]=\delta(x)\in\delta({{}_{s}X})\subseteq{{}_{s}E}.$
@@
$[a,y]=\delta(x)=(\delta\circ\alpha_{s})(z)=(\beta_{s}\circ\delta)(z)\in E_{s}.$
In both cases, we have a contradiction with $[a,y]\in E_{0}$. Therefore, $E_{0}\cap\delta(X)\subseteq\delta(X_{0})$. Now, chose an element $[a,y]\in E\setminus\delta(X)$. Then, we can fix a representative $(s,x)\simeq(a,y)$ such that $s\in S$. Since $(s,x)\in D$ and $s\in S$, it follows from the definition of $D$ that
@@
$[\delta,x]\in{{}_{s}E}\quad\text{and}\quad\beta_{s}([\delta,x])=[s,x]\in E_{s}.$
On the other hand, if $x\in X_{t}$, for some $t\in S^{s}$, then there is $z\in X$ such that $x=\alpha_{t}(z)$. Therefore
@@
That is, $E\setminus\delta(X)\subseteq E_{1}$. Therefore $E_{0}=E_{0}\cap\delta(X)\subseteq\delta(X_{0})\subseteq E_{0}$. Since a partial action $(\alpha,X)$ is non-degenerate if and only if $X_{0}=\emptyset$, we conclude that that $(\alpha,X)$ is non-degenerate if and only if $\emptyset=\delta(X_{0})=E_{0}$ if and only if $(\delta,\beta,E)$ is non-degenerate. ∎
@@ 4. Application in semigroups and categories
In this subsection we compare our definition of partial action when restricted to a semigroup with the one presented in [17]. Furthermore, we show that the universal globalization constructed in Section 3 coincides with the tensor product globalization when the partial semigroup action is non-degenerate.
In this subsection, $\mathcal{S}$ will denote a semigroup (that is, a semigroupoid such that $\mathcal{S}^{(2)}=\mathcal{S}\times\mathcal{S}$).
@@
[17, Definition 2.3] A strong (left) partial semigroup action of $\mathcal{S}$ on $X$ is a pair ($(\{{{}_{s}X}\}_{s\in\mathcal{S}},\{\alpha_{s}\}_{s\in\mathcal{S}})$ consisting of a collection $\{_{s}X\}_{s\in\mathcal{S}}$ of subsets of $X$ and a collection $\{\alpha_{s}\}_{s\in\mathcal{S}}$ of functions $\alpha_{g}\colon{{}_{g}X}\to X$ such that:
@@
The strong partial actions described in [17] are the type of partial actions that eventually can be globalizable [17, Proposition 2.9]. In [17, Lemma 2.4] it has been shown that strong partial monoid actions (in the sense of [15, Definition 2.4]) are precisely the strong and unitary partial semigroup actions of a monoid. We also observe that global semigroup actions are always non-degenerate, whereas partial actions are not necessarily so.
@@
Since $\mathcal{S}^{(2)}=\mathcal{S}\times\mathcal{S}$, condition (S) is equivalent to (P1) and (P2), so (i) is proved.
@@
Now we relate our definition of (universal) globalization to the one presented for semigroups.
The notion of globalization presented in [17, Definition 2.8] is analogous to ours. That is, a globalization of a partial semigroup action $(\alpha,X)$ is a triple $(i,\gamma,Y)$ where $(\gamma,Y)$ is a global semigroup action and $i\colon(\alpha,X)\to(\gamma,Y)$ is an embedding. On the other hand, the term universal globalization was not adopted in [17], as two different globalizations of a partial semigroup action were presented, each one being universal in some sense. The globalization we are interested in is called tensor product globalization. It is constructed as follows.
@@
($\ast$) $(s,x)\sim(t,y)\text{ if there is }u\in\mathcal{S}^{1}\text{ such that }s=ut,x\in{{}_{u}X}\text{ and }\alpha_{u}(x)=y.$
Let $\mathcal{S}^{1}\otimes X$ be the quotient set $(\mathcal{S}^{1}\times X)/\approx$ and define $\gamma_{s}(t\otimes x)=st\otimes x$, for every $t\otimes x\in\mathcal{S}^{1}\otimes X$. This is a global action of $\mathcal{S}$ on $\mathcal{S}^{1}\otimes X$ and the function $i\colon X\to\mathcal{S}^{1}\otimes X$, defined as $i(x)=1\otimes x$, is an embedding. This triple $(i,\gamma,\mathcal{S}^{1}\otimes\delta)$ is the tensor product globalization of $(\alpha,X)$.
It has been shown in [17, Theorem 3.5] that $(i,\gamma,\mathcal{S}^{1}\otimes X)$ is an initial object in the category of globalizations of $(\alpha,X)$, denoted by $\mathcal{G}(\alpha,X)$. The category $\mathcal{G}(\alpha,X)$ has as objects the embeddings $\delta\colon(\alpha,X)\to(\beta,Z)$, where $(\beta,Z)$ is a global semigroup action, and the morphisms between the objects $\delta\colon(\alpha,X)\to(\beta,Z)$ and $\varphi\colon(\alpha,X)\to(\theta,W)$ are the morphisms $\Phi\colon(\beta,Z)\to(\theta,W)$ such that $\Phi\circ\delta=\varphi$.
@@
    Let $\mathcal{A}^{\ast}(\mathcal{S})$ be the full subcategory of $\mathcal{A}(\mathcal{S})$ whose objects are the non-degenerated global actions of $\mathcal{S}$, and $I^{\ast}\colon\mathcal{A}^{\ast}(\mathcal{S})\to\mathcal{A}_{p}(\mathcal{S})$ be the inclusion functor. Then $((\alpha,X)\to I^{\ast})$ is a proper full subcategory of $((\alpha,X)\to I)$ (recall the definition in Proposition 2.23).
@@
    The category of globalizations $\mathcal{G}(\alpha,X)$ is a proper full subcategory of $((\alpha,X)\to I^{\ast})$. In the category $((\alpha,X)\to I^{\ast})$ the objects consist of all morphisms of partial actions $(\alpha,X)\to(\beta,Y)$, whereas $\mathcal{G}(\alpha,X)$ restricts its objects to embeddings only.
@@
    The proof given in [17, Theorem 3.5] does not use the hypothesis that an object $\delta\in\mathcal{G}(\alpha,X)$ is an embedding. Thus, the tensor product globalization $(i,\gamma,\mathcal{S}^{1}\otimes X)$ also provides an initial object for the category $((\alpha,X)\to I^{\ast})$.
Due to Proposition 3.8, the universal globalization $(\delta,\beta,E)$ of a partial action $(\alpha,X)$, given by Theorem 3.7, is non-degenerate if and only if $(\alpha,X)$ is non-degenerate. On the other hand, any global semigroup action (in the sense of Kudryavtseva and Laan) is non-degenerate. In particular, the tensor product globalization $(i,\gamma,\mathcal{S}^{1}\otimes X)$ is non-degenerate. In the next result, we relate the globalizations $(\delta,\beta,E)$ and $(i,\gamma,\mathcal{S}^{1}\otimes X)$.
@@
Suppose that $(\alpha,X)$ is non-degenerate. From Proposition 3.8, it follows that $(\delta,\beta,Y)$ is non-degenerate and, hence, $(\delta,\beta,Y)\in((\alpha,X)\to I^{\ast})$. By Proposition 2.23(i), $\delta$ is an initial object in the category $((\alpha,X)\to I)$. Since $((\alpha,X)\to I^{\ast})$ is a full subcategory of $((\alpha,X)\to I)$, it follows that $\delta$ is an initial object in the category $((\alpha,X)\to I^{\ast})$.
On the other hand, $i\colon(\alpha,X)\to(\gamma,\mathcal{S}^{1}\otimes X)$ is an initial object in the category $(\alpha(X)\to I^{\ast})$. Since initial objects are unique, up to isomorphism, there is an isomorphism $\Phi\colon\delta\to i$. By definition, $\Phi$ is an isomorphism of partial actions $(\beta,E)\to(\gamma,\mathcal{S}^{1}\otimes X)$.

### v2

@@ 3. The universal globalization
(i) We first verify (G1). If $s,t\in S$ are such that $S^{s}=S^{t}\neq\emptyset$ and $a,x]\in{{}_{s}E}$, then $[a,x]=[r,y]$ for some $r\in S^{s}$, or $a,x]=[\delta,y]$ for some $y\in{{}_{\rho_{s}}X}$. In the first case we have $r\in S^{s}=S^{t}$ and, hence, $a,x]\in{{}_{t}E}$. In the second, $S^{s}=S^{t}\neq\emptyset$ implies $\rho_{s}=\rho_{t}$, hence $y\in{{}_{\rho_{t}}X}$ and $a,x]=[\delta,y]\in{{}_{t}E}$. This shows ${{}_{s}E}\subseteq{{}_{t}E}$. Since $S^{s}=S^{t}\neq\emptyset$ is a symmetric condition, the same argument proves ${{}_{s}E}={{}_{t}E}$. For the remaining of the proof for (i) we uses Lemma 2.10(ii). That is, for $(s,t)\in S^{(2)}$ we must show ${{}_{t}E}={{}_{st}E}$, $\beta_{t}({{}_{t}E})\subseteq{{}_{s}E}$ and $\beta_{s}\circ\beta_{t}=\beta_{st}$.
If $a,x]\in{{}_{t}E}$, then $\beta_{t}([a,x])=[tr,y]$ for some $r\in S^{t}$ or $\beta_{t}([a,x])=[t,y]$. Since $t,tr\in S^{s}$, in both cases we have $\beta_{t}([a,x])\in{{}_{s}E}$. Thus $E_{t}\subseteq{{}_{s}E}$. Since $(s,t)\in S^{(2)}$ implies $S^{t}=S^{st}$ and, by Lemma 3.2(i), $\rho_{t}=\rho_{st}$, we have
@@
(ii) We first verify that $(\iota,\beta,E)$ is a globalization for $(\alpha,X)$. Let $x\in{{}_{s}X}$ and notice that ${{}_{s}X}\subseteq{{}_{\rho_{s}}X}$. Then $(\delta,x)\in{{}_{s}D}$ and, therefore, $\iota(x)=[\delta,x]\in{{}_{s}E}$. Furthermore, for such $x$ we have
@@
Thus, $\iota$ is a morphism of partial actions. It follows from Lemma 3.4(v) that $\iota$ is injective. Therefore, to $\iota$ to be an embedding, it remains to prove that ${{}_{s}X}=\{x\in X\colon\iota(x)\in{{}_{s}E},\beta_{s}(\iota(x))\in\iota(X)\}$. Suppose that $x\in X$ is such that $\iota(x)\in{{}_{s}E}$ and $\beta_{s}(\iota(x))\in\iota(X)$. From $\iota(x)\in{{}_{s}E}$ we have two possible cases:
@@
    $\delta,x]=[\delta,y]$ for some $y\in{{}_{\rho_{s}}X}$. In this case, Lemma 3.4(v) implies $x=y$. Therefore $\beta_{s}(\iota(x))=[s,x]$. Now, if $s,x]\in\iota(X)$, then there is $z\in X$ such that $s,x]=[\delta,z]$. As any chain from $(\delta,z)$ to $(s,x)$ must use relation (1), it follows from Lemma 3.4(ii) that $x\in{{}_{s}X}$.
@@
    $\delta,x]=[t,y]$, for some $t\in S^{s}$. Similar to the previous case, we obtain from Lemma 3.4(ii) that $y\in{{}_{t}X}$ and $\alpha_{t}(y)=x$. Now, if $\beta_{s}(\iota(x))=[st,y]\in\iota(X)$, then there is $z\in X$ such that $st,y]=[\delta,z]$. Again, it must be $y\in{{}_{st}X}$. Therefore
@@
This shows that $\{x\in X\colon\iota(x)\in{{}_{s}E},\beta_{s}(\iota(x))\in\iota(X)\}\subseteq{{}_{s}X}$. The reverse inclusion follows from $\iota$ being a morphism of partial actions. This concludes that $(\iota,\beta,E)$ is a globalization for $(\alpha,X)$.
Now we verify that $(\iota,\beta,E)$ is a universal globalization. To prove this, let $(\theta,Y)$ be a global action and $\varphi\colon(\alpha,X)\to(\theta,Y)$ be a morphism. Define $\Phi\colon D\to Y$ as
@@
For $\Phi$ to be well defined, we need to verify that $\varphi(x)\in{{}_{s}Y}$ whenever $(s,x)\in D$. In fact, if $(s,x)\in D$, then $x\in{{}_{\rho_{s}}X}$ and, hence, we have two possible cases:
@@
    $x\in{{}_{t}X}$ for some $t\in S$ such that $sRt$. In this case, since $\varphi$ is a morphism of partial actions, we have $\varphi(x)\in{{}_{t}Y}$. As $(\theta,Y)$ is a global action, it follows from Lemma 3.2(iii) that $\varphi(x)\in{{}_{t}Y}={{}_{s}Y}$.
@@
Hence, the function $\Phi$ is well defined. Furthermore, $\Phi$ is compatible with the relation $\sim$ on $D$. For if $(s,x)\sim(\delta,y)$, then $x\in{{}_{s}X}$ and $y=\alpha_{s}(x)$. Therefore,
@@
While if $(s,x)\sim(t,y)$, then there is $u\in S^{t}$ such that $s=tu$, $x\in{{}_{u}X}$ and $\alpha_{u}(x)=y$. Therefore,
@@
The case $t=su$ is analogous. As $\Phi$ is compatible with $\sim$ and, consequently, with $\approx$, it induces a function $\Phi\colon E\to Y$, given by $\Phi([a,x])=\Phi(a,x)$. The function $\Phi$ is a morphism of partial actions and satisfy $\Phi\circ\iota=\varphi$. For if $t,x]\in{{}_{s}E}$, where $t\in S^{s}$, then
@@
while if $\delta,x]\in{{}_{s}E}$, then
$(\Phi\circ\beta_{s})([\delta,x])=(\theta_{s}\circ\varphi)(x)=(\theta_{s}\circ\Phi)([\delta,x]).$
Thus, $\Phi$ is a morphism of partial action. For every $x\in X$, we have $(\Phi\circ\iota)(x)=\Phi([\delta,x])=\phi(x)$. For the uniqueness of $\Phi$, suppose that $\Psi\colon E\to Y$ is a morphism such that $\Psi\circ\iota=\varphi=\Phi\circ\iota$. For each $\delta,x]\in E$, we have $\Psi([\delta,x])=(\Psi\circ\iota)(x)=(\Phi\circ\iota)(x)=\Phi([\delta,x])$. Now, for each $s,x]\in E$, we have $x\in{{}_{\rho_{s}}X}$. Therefore $s,x]=\beta_{s}([\delta,x])$ and, hence,
@@
That is, $\Psi\equiv\Phi$. Therefore $(\iota,\beta,E)$ is a universal globalization for $(\alpha,X)$. ∎
To end this section, we will prove that the universal globalization of a partial semigroupoid action is non-degenerate if and only if the partial action is non-degenerate. Let $(\alpha,X)$ be a partial semigroupoid action. We denote
@@
That is, $X_{1}$ and $X_{0}$ are, respectively, the non-degenerate and the degenerate parts of $X$. With this notation, $(\alpha,X)$ is non-degenerate if and only if $X_{1}=X$ if and only if $X_{0}=\emptyset$.
@@
Let $(\alpha,X)$ be a partial semigroupoid action of $S$, and let $(\iota,\beta,E)$ be its universal globalization, as stated in Theorem 3.7. Then $\iota(X_{0})=E_{0}$. Consequently, $(\alpha,X)$ is non-degenerate if and only if $(\beta,E)$ is non-degenerate.
@@
First, we prove the inclusion $\iota(X_{0})\subseteq E_{0}$. Let $x\in X_{0}$ and $\iota(x)=[\delta,x]\in E$. Suppose that $\iota(x)\in{{}_{s}E}$, for some $s\in S$, and recall that ${{}_{s}E}={{}_{s}D}/\approx$, where
@@
Since $x\in X_{0}$, it follows that $x\notin{{}_{\rho_{s}}X}$. Therefore, $(\delta,x)\approx(t,y)$ for some $t\in S^{s}$. This implies that there exists a chain of the form
@@
for some $r\in S$. However, $(\delta,x)\sim(r,z)$ means that $x=\alpha_{r}(z)\in X_{r}\subseteq X_{1}$, a contradiction. Now, suppose that $\iota(x)\in E_{s}$, for some $s\in S$. Then, it must be
$\delta,x]=\beta_{s}([\delta,y])=[s,y]\quad\text{or}\quad[\delta,x]=\beta_{s}([t,y])=[st,y].$
Both cases lead to $(\delta,x)\approx(r,y)$, for some $r\in S$, which, again, result in a contradiction. Therefore,
$\iota(x)\notin\cup_{s\in S}({{}_{s}E}\cup E_{s})=E_{1}\implies\iota(x)\in E\setminus E_{1}=E_{0}.$
This concludes that $\iota(X_{0})\subseteq E_{0}$.
To prove the inclusion $E_{0}\subseteq\iota(X_{0})$, we will first consider the case $a,y]\in E_{0}\cap\iota(X)$. That is, $a,y]=[\delta,x]$, for some $x\in X$. Suppose that $x\in X_{1}$. Then there is $s\in S$ such that $x\in{{}_{s}X}\cup X_{s}$. Since $\iota\colon(\alpha,X)\to(\beta,E)$ is a morphism of partial actions, if $x\in{{}_{s}X}$, then
$a,y]=\iota(x)\in\iota({{}_{s}X})\subseteq{{}_{s}E}.$
@@
$a,y]=\iota(x)=(\iota\circ\alpha_{s})(z)=(\beta_{s}\circ\iota)(z)\in E_{s}.$
In both cases, we have a contradiction with $a,y]\in E_{0}$. Therefore, $E_{0}\cap\iota(X)\subseteq\iota(X_{0})$. Now, chose an element $a,y]\in E\setminus\iota(X)$. Then, we can fix a representative $(s,x)\approx(a,y)$ such that $s\in S$. Since $(s,x)\in D$ and $s\in S$, it follows from the definition of $D$ that
@@
$\delta,x]\in{{}_{s}E}\quad\text{and}\quad\beta_{s}([\delta,x])=[s,x]\in E_{s}.$
On the other hand, if $x\in X_{t}$, for some $t\in S^{s}$, then there is $z\in X$ such that $x=\alpha_{t}(z)$. Therefore,
@@
That is, $E\setminus\iota(X)\subseteq E_{1}$. Hence, $E_{0}=E_{0}\cap\iota(X)\subseteq\iota(X_{0})\subseteq E_{0}$. Since a partial action $(\alpha,X)$ is non-degenerate if and only if $X_{0}=\emptyset$, we conclude that that $(\alpha,X)$ is non-degenerate if and only if $\emptyset=\iota(X_{0})=E_{0}$ if and only if $(\beta,E)$ is non-degenerate. ∎
@@ 4. Application in semigroups and categories
In this subsection, we compare our definition of partial action when restricted to a semigroup with the one from [19]. Furthermore, we show that the universal globalization constructed in Section 3 coincides with the tensor product globalization when the partial semigroup action is non-degenerate.
Throughout this subsection, $\mathcal{S}$ will denote a semigroup (that is, a semigroupoid such that $\mathcal{S}^{(2)}=\mathcal{S}\times\mathcal{S}$).
@@
[19, Definition 2.3] A strong (left) partial semigroup action of $\mathcal{S}$ on $X$ is a pair $(\{{{}_{s}X}\}_{s\in\mathcal{S}},\{\alpha_{s}\}_{s\in\mathcal{S}})$ consisting of a collection $\{_{s}X\}_{s\in\mathcal{S}}$ of subsets of $X$ and a collection $\{\alpha_{s}\}_{s\in\mathcal{S}}$ of functions $\alpha_{s}\colon{{}_{s}X}\to X$ such that:
@@
The strong partial actions described in [19] are the type of partial actions that eventually can be globalizable [19, Proposition 2.9]. In [19, Lemma 2.4] it has been shown that strong partial monoid actions (in the sense of [16, Definition 2.4]) are precisely the strong and unitary partial semigroup actions of a monoid. We also observe that global semigroup actions are always non-degenerate, whereas partial actions are not necessarily so.
@@
Since $\mathcal{S}^{(2)}=\mathcal{S}\times\mathcal{S}$, condition (S) is equivalent to (P1) and (P2). This proves (i).
@@
Now, we relate our definition of (universal) globalization to the one presented for semigroups.
Our notion of globalization is analogous to the respective notion from [19, Definition 2.8]. That is, a globalization of a partial semigroup action $(\alpha,X)$ is a triple $(i,\gamma,Y)$ where $(\gamma,Y)$ is a global semigroup action and $i\colon(\alpha,X)\to(\gamma,Y)$ is an embedding. On the other hand, the term universal globalization was not adopted in [19], as two different globalizations of a partial semigroup action were considered, each one being universal in a certain sense. The globalization we are interested in is called the tensor product globalization. It is constructed as follows.
@@
[…]
Let $\mathcal{S}^{1}\otimes X$ be the quotient set $(\mathcal{S}^{1}\times X)/\approx$ and define $\gamma_{s}(t\otimes x)=st\otimes x$, for every $t\otimes x\in\mathcal{S}^{1}\otimes X$. This is a global action of $\mathcal{S}$ on $\mathcal{S}^{1}\otimes X$ and the function $i\colon X\to\mathcal{S}^{1}\otimes X$, defined as $i(x)=1\otimes x$, is an embedding. This triple $(i,\gamma,\mathcal{S}^{1}\otimes X)$ is the tensor product globalization of $(\alpha,X)$.
It has been shown in [19, Theorem 3.5] that $(i,\gamma,\mathcal{S}^{1}\otimes X)$ is an initial object in the category of globalizations of $(\alpha,X)$, denoted by $\mathcal{G}(\alpha,X)$. The category $\mathcal{G}(\alpha,X)$ has as objects the embeddings $\iota\colon(\alpha,X)\to(\beta,Z)$, where $(\beta,Z)$ is a global semigroup action, and the morphisms between the objects $\iota\colon(\alpha,X)\to(\beta,Z)$ and $\varphi\colon(\alpha,X)\to(\theta,W)$ are the morphisms $\Phi\colon(\beta,Z)\to(\theta,W)$ such that $\Phi\circ\iota=\varphi$.
@@
    Let $\mathcal{A}^{\ast}(\mathcal{S})$ be the full subcategory of $\mathcal{A}(\mathcal{S})$ whose objects are the non-degenerated global actions of $\mathcal{S}$, and $I^{\ast}\colon\mathcal{A}^{\ast}(\mathcal{S})\to\mathcal{A}_{p}(\mathcal{S})$ be the inclusion functor. Then $((\alpha,X)\downarrow I^{\ast})$ is a proper full subcategory of $((\alpha,X)\downarrow I)$ (recall the definition in Proposition 2.23).
@@
    The category of globalizations $\mathcal{G}(\alpha,X)$ is a proper full subcategory of $((\alpha,X)\downarrow I^{\ast})$. In the category $((\alpha,X)\downarrow I^{\ast})$ the objects consist of all morphisms of partial actions $(\alpha,X)\to(\beta,Y)$, whereas $\mathcal{G}(\alpha,X)$ restricts its objects to embeddings only.
@@
    The proof given in [19, Theorem 3.5] does not use the hypothesis that an object $\iota\in\mathcal{G}(\alpha,X)$ is an embedding. Thus, the tensor product globalization $(i,\gamma,\mathcal{S}^{1}\otimes X)$ also provides an initial object for the category $((\alpha,X)\downarrow I^{\ast})$.
Due to Proposition 3.8, the universal globalization $(\iota,\beta,E)$ of a partial action $(\alpha,X)$, given by Theorem 3.7, is non-degenerate if and only if $(\alpha,X)$ is non-degenerate. On the other hand, any global semigroup action (in the sense of Kudryavtseva and Laan) is non-degenerate. In particular, the tensor product globalization $(i,\gamma,\mathcal{S}^{1}\otimes X)$ is non-degenerate. In the next result, we relate the globalizations $(\iota,\beta,E)$ and $(i,\gamma,\mathcal{S}^{1}\otimes X)$.
@@
Suppose that $(\alpha,X)$ is non-degenerate. From Proposition 3.8, it follows that $(\iota,\beta,Y)$ is non-degenerate and, hence, $(\iota,\beta,Y)\in((\alpha,X)\downarrow I^{\ast})$. By Proposition 2.23(i), $\iota$ is an initial object in the category $((\alpha,X)\downarrow I)$. Since $((\alpha,X)\downarrow I^{\ast})$ is a full subcategory of $((\alpha,X)\downarrow I)$, it follows that $\iota$ is an initial object in the category $((\alpha,X)\downarrow I^{\ast})$.
On the other hand, $i\colon(\alpha,X)\to(\gamma,\mathcal{S}^{1}\otimes X)$ is an initial object in $(\alpha(X)\downarrow I^{\ast})$. Since initial objects are unique, up to isomorphism, there is an isomorphism $\Phi\colon\iota\to i$. By definition, $\Phi$ is an isomorphism of partial actions $(\beta,E)\to(\gamma,\mathcal{S}^{1}\otimes X)$.